# Imperfect dark matter with higher derivatives

### v1

@@ 4.2 Landau frame
In general, $q^{\mu}\neq 0$ in our setup. In the comoving frame, we found $q^{\mu}\propto\big(C_{\phi X}\,X+2C_{XX}\,Y\big)\,\chi^{\mu}$, so it vanishes only if $\chi^{\mu}=0$ (e.g. on FLRW) or if the prefactor happens to vanish. Alternatively, one may choose the Landau (energy) frame, defined by setting the energy flux to zero by construction. In this frame the four-velocity is the timelike eigenvector of $T^{\mu}{}_{\nu}$,
$T^{\mu}{}_{\nu}\,u^{\nu}=-\rho\,u^{\mu}\,.$ (4.14)
@@
From Eq. (4.16) we see that if $\Delta=0$ then $p=0$ and the rank of $T^{\mu}{}_{\nu}$ drops from $2$ to $1$, i.e., the stress–energy reduces to that of dust. This occurs in two distinct situations. First, if $X_{\mu}=\eta\,\phi_{\mu}$ (for some scalar $\eta$), then by the definitions (2.19) one has $XZ-Y^{2}=0$, and $T_{\mu\nu}=-2\lambda\!\left(C_{\phi\phi}+\eta\,C_{\phi X}+\eta^{2}C_{XX}\right)\phi_{\mu}\phi_{\nu}$ which is rank one, i.e., proportional to the square of the single timelike one-form that is proportional to $\phi_{\mu}$. Secondly, if $C_{\phi\phi}C_{XX}-\tfrac{1}{4}C_{\phi X}^{2}=0$, $T_{\mu\nu}$ is proportional to the squared of the single timelike 1-form $\phi_{\mu}+\sqrt{\tfrac{C_{XX}}{C_{\phi\phi}}}X_{\mu}$. The usual Schutz (dust) case with action (1.2) is a special subset of the latter case with $C_{\phi X}=C_{XX}=0$, yielding $T_{\mu\nu}=2\lambda\,\phi_{\mu}\phi_{\nu}$ and $\rho=2\lambda\,C_{\phi\phi}X=2\lambda$ where we used $C_{\phi\phi}=f=1/X$ (so that $Xf=1$).
@@ 5 Avoiding caustics
In this section we briefly review the standard focusing of geodesic dust, and then show that in the higher-derivative model with energy–momentum tensor (3.4) the additional terms generically induce non-geodesic acceleration and can source vorticity, thereby regulating the flow and helping to prevent caustic formation.
@@ 5.2 Higher derivatives: non-vanishing acceleration and vorticity
In the dust case (1.1), the acceleration and vorticity vanish, $a^{\mu}=0$ and $\omega_{\mu\nu}=0$, so the Raychaudhuri equation (5.6) integrates to a finite focusing time (5.12) for the formation of caustic singularities. In contrast, for the higher-derivative energy–momentum tensor (3.4) we show that both $a^{\mu}$ and $\omega_{\mu\nu}$ are generically nonzero.
@@
where $D_{\mu}$ is the spatial covariant derivative defined in (5.5). Since $a^{\mu}$ is spatial ($u_{\alpha}a^{\alpha}=0$), one has $a^{\alpha}a_{\alpha}\geq 0$. Thus the last term in (5.14) contributes positively on the right-hand side of (5.6), opposing geodesic focusing and helping to prevent caustic formation. The sign of the first term, $D_{\alpha}a^{\alpha}$, is not fixed in general. To get better understanding of $D_{\alpha}a^{\alpha}$, let us obtain an explicit expression for the acceleration. In order to do so, we work in the Landau frame ($q^{\mu}=0$). Projecting $\nabla_{\alpha}T^{\mu\alpha}=0$ orthogonally to $u^{\mu}$ yields the relativistic Euler equation
@@
The vector $F_{\mu}$ is the relativistic analogue of a force density, so spatial gradients of $p$ and $\pi^{\mu\nu}$ generically imply $a_{\mu}\neq 0$. An analogous but more involved expression holds in the scalar comoving frame (where $q^{\mu}\neq 0$). From (5.14) and (5.15), the sign of $D_{\alpha}a^{\alpha}$ depends on spatial gradients of $\rho+p$, $p$, and $\pi^{\mu\nu}$ and can take either sign.
@@
The idea that higher-derivative structures can prevent caustic formation has been appeared in other modified gravity scenarios. In projectable Hořava–Lifshitz gravity, it is argued that higher-curvature terms can prevent formation of caustics singularities [58]. Moreover, several works make explicit the close connection between mimetic gravity and the infrared limit of projectable Hořava–Lifshitz gravity, suggesting a common mechanism whereby the extra mode sources non-geodesic flow [59, 60, 61].
@@ 6 Summary and conclusions
Nevertheless, the mimetic construction provides a systematic approach to the dust action by looking at the singular limit of a conformal transformation of the metric in the gravitational sector. In this work we followed that approach to incorporate higher-derivative effects in a systematic way. In particular, consistency of the underlying disformal map restricts the allowed higher-derivative structures, leading to the generalized action (3.1). Unlike earlier approaches, where higher-derivative operators were added directly to the original mimetic Lagrangian, here all higher-derivative corrections are encoded solely in the (generalized) mimetic constraint (3.2). This structure automatically renders the theory free of Ostrogradsky-like ghost degrees of freedom, as the higher derivatives of the scalar field can be re-expressed in terms of lower derivatives. The resulting energy–momentum tensor describes an imperfect fluid and reduces to dust when the higher-derivative terms are switched off. We showed that, on homogeneous cosmological backgrounds the theory behaves as pressureless dark matter, whereas in the presence of inhomogeneities the higher-derivative terms generate nonzero acceleration (non-geodesic flow) and vorticity, thereby helping to prevent caustic formation.
@@ Acknowledgements
This work was supported by IBS under the project code IBS-R018-D3.

### v2

@@ 4.2 Landau frame
In general, $q^{\mu}\neq 0$ in our setup. In the comoving frame we found $q^{\mu}\propto\big(C_{\phi X}\,X+2C_{XX}\,Y\big)\,V^{\mu}$, so it vanishes only if $V^{\mu}=0$ (for instance on an FLRW background) or if the prefactor happens to vanish, as in the subclass $f({\cal C})$ with $\cal C$} defined in Eq. (3.7). Alternatively, one may choose the Landau (energy) frame, defined by setting $q^{\mu}=0$ by construction. In this frame the four-velocity is the timelike eigenvector of $T^{\mu}{}_{\nu}$,
$T^{\mu}{}_{\alpha}\,u^{\alpha}=-\rho\,u^{\mu}\,.$ (4.18)
@@
From Eq. (4.20) we see that if $\Delta=0$ then $p=0$ and the rank of $T^{\mu}{}_{\nu}$ drops from $2$ to $1$, i.e., the energy-momentum tensor reduces to that of dust. This occurs in two distinct situations. First, if $X_{\mu}=\eta\,\phi_{\mu}$ (for some scalar $\eta$), then by the definitions (2.19) one has $XZ-Y^{2}=0$, and $T_{\mu\nu}=-2\lambda\!\left(C_{\phi\phi}+\eta\,C_{\phi X}+\eta^{2}C_{XX}\right)\phi_{\mu}\phi_{\nu}$ which is rank one, i.e., proportional to the square of the single timelike one-form that is proportional to $\phi_{\mu}$. Secondly, if $C_{\phi\phi}C_{XX}-\tfrac{1}{4}C_{\phi X}^{2}=0$, $T_{\mu\nu}$ is proportional to the squared of the single timelike 1-form $\phi_{\mu}+\sqrt{\tfrac{C_{XX}}{C_{\phi\phi}}}X_{\mu}$. The usual Schutz (dust) case with action (1.2) is a special subset of the latter case with $C_{\phi X}=C_{XX}=0$, yielding $T_{\mu\nu}=2\lambda\,\phi_{\mu}\phi_{\nu}$ and $\rho=2\lambda\,C_{\phi\phi}X=2\lambda$ where we used $C_{\phi\phi}=f=1/X$ (so that $Xf=1$).
@@ 5 Avoiding caustics
In this section we briefly review the standard focusing of geodesic dust, and then show that in the higher-derivative model with energy-momentum tensor (3.4) the additional terms generically induce non-geodesic acceleration and can source vorticity, thereby regulating the flow and helping to prevent caustic formation.
@@ 5.2 Higher derivatives: non-vanishing acceleration and vorticity
In the dust case (1.1), the acceleration and vorticity vanish, $a^{\mu}=0$ and $\omega_{\mu\nu}=0$, so the Raychaudhuri equation (5.6) integrates to a finite focusing time (5.12) for the formation of caustic singularities. In contrast, for the higher-derivative energy-momentum tensor (3.4) we show that both $a^{\mu}$ and $\omega_{\mu\nu}$ are generically nonzero.
@@
where $D_{\mu}$ is the spatial covariant derivative defined in (5.5). Since $a^{\mu}$ is spatial ($u_{\alpha}a^{\alpha}=0$), one has $a^{\alpha}a_{\alpha}\geq 0$. Thus the last term in (5.14) contributes positively on the right-hand side of (5.6), opposing geodesic focusing and helping to prevent caustic formation. The sign of the first term, $D_{\alpha}a^{\alpha}$, is not fixed in general. To get better understanding of $D_{\alpha}a^{\alpha}$, let us obtain an explicit expression for the acceleration. It is useful to start with the scalar comoving frame $u^{\mu}\propto\phi^{\mu}$, where one has $a_{\mu}=-V_{\mu}/(2X)$. Taking a spatial divergence gives
$D_{\alpha}a^{\alpha}=-\frac{1}{2X}\,D_{\alpha}V^{\alpha}+\frac{1}{2X^{2}}\,V^{\alpha}D_{\alpha}X,\qquad\mbox{comoving frame}$ (5.15)
so while $a_{\alpha}a^{\alpha}\geq 0$ is sign-definite, the sign of $D_{\alpha}a^{\alpha}$ depends on the local spatial profile of $X$ (and thus on inhomogeneities). For example, in regimes where spatial gradients are dominated by the scalar $X$ itself so that $V_{\mu}\simeq D_{\mu}X$ and $D_{\alpha}V^{\alpha}\simeq D^{2}X$, one has $D_{\alpha}a^{\alpha}\simeq-(2X)^{-1}D^{2}X+\cdots$; since $X<0$ on the timelike branch, a local condition such as $D^{2}X>0$ then tends to give $D_{\alpha}a^{\alpha}>0$ and hence contributes to defocusing in the Raychaudhuri equation. More generally, however, $D_{\alpha}a^{\alpha}$ can have either sign because both terms in (5.15) contribute. In the Landau frame ($q^{\mu}=0$), projecting $\nabla_{\alpha}T^{\mu\alpha}=0$ orthogonally to $u^{\mu}$ yields the relativistic Euler equation
@@
The vector $F_{\mu}$ is the relativistic analogue of a force density, so spatial gradients of $p$ and $\pi^{\mu\nu}$ generically imply $a_{\mu}\neq 0$.
@@
It should be stressed that the above criterion is meant as a structural indication rather than an explicit proof of caustic avoidance in a given configuration: establishing whether (5.19) is actually satisfied along the flow for representative inhomogeneous solutions would require solving the coupled system (analytically in symmetry-reduced setups or numerically), which is beyond the scope of the present work. Nevertheless, an important consistency aspect of the subclass $f({\cal C})$ identified in Sec. 3 is that, in unitary gauge, the higher-derivative contributions enter through spatial derivatives only. As a result, the evolution equations in this sector do not acquire higher time derivatives, and the initial-value formulation is not expected to be obstructed by Ostrogradsky-type pathologies, while the new spatial-gradient terms can become relevant precisely in regimes where inhomogeneities grow and the standard dust congruence would otherwise undergo geodesic focusing. This provides a concrete setting in which the mechanism suggested by (5.19) can operate, and motivates a dedicated study of explicit solutions and well-posedness properties as a separate follow-up.
The idea that higher-derivative structures can prevent caustic formation has been appeared in other modified gravity scenarios. In projectable Hořava-Lifshitz gravity, it is argued that higher-curvature terms can prevent formation of caustics singularities [66]. Moreover, several works make explicit the close connection between mimetic gravity and the infrared limit of projectable Hořava-Lifshitz gravity, suggesting a common mechanism whereby the extra mode sources non-geodesic flow [67, 68, 69, 70].
@@ 6 Summary and conclusions
Nevertheless, the mimetic construction provides a systematic approach to the dust action by looking at the singular limit of a conformal transformation of the metric in the gravitational sector. In this work we followed that approach to incorporate higher-derivative effects in a systematic way. In particular, consistency of the underlying disformal map restricts the allowed higher-derivative structures, leading to the generalized action (3.1). Unlike earlier approaches, where higher-derivative operators were added directly to the original mimetic Lagrangian, here all higher-derivative corrections are encoded solely in the (generalized) mimetic constraint (3.2). In general, establishing the absence of Ostrogradsky-like ghost degrees of freedom for the full theory (3.1) requires a complete Hamiltonian analysis. On the other hand, there exists a particularly simple and well-controlled subclass, obtained by restricting to $f({\cal C})$ with $\cal C$} defined in (3.7), for which the potentially dangerous higher-time-derivative structure is absent already at the level of the operator basis (and in unitary gauge the higher-derivative contributions enter through spatial derivatives only). The resulting energy-momentum tensor describes an imperfect fluid and reduces to dust when the higher-derivative terms are switched off. We showed that, on homogeneous cosmological backgrounds the theory behaves as pressureless dark matter, whereas in the presence of inhomogeneities the higher-derivative terms generate nonzero acceleration (non-geodesic flow) and vorticity, thereby helping to prevent caustic formation.
@@ Acknowledgements
I am grateful to Karim Noui and Pavel Petrov for helpful discussions. This work was supported by IBS under the project code IBS-R018-D3.